40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 16 19:34:18 2021, Koji, Update, General, HVAC HVAC_Power.jpegSouth_Arm.jpegSouth_End_Tenperature.jpegVertex.jpegEast_Arm.jpeg
    Reply  Wed Jul 14 16:19:44 2021, gautam, Update, General, Brrr rmTemp.pdf
       Reply  Wed Jul 14 19:21:44 2021, Koji, Update, General, Brrr 
Message ID: 16245     Entry time: Wed Jul 14 16:19:44 2021     In reply to: 16206     Reply to this: 16246
Author: gautam 
Type: Update 
Category: General 
Subject: Brrr 

Since the repair work, the temperature is significantly cooler. Surprisingly, even at the vertex (to be more specific, inside the PSL enclosure, which for the time being is the only place where we have a logged temperature sensor, but this is not attributable to any change in the HEPA speed), the temperature is a good 3 deg C cooler than it was before the HVAC work (even though Koji's wind vane suggest the vents at the vertex were working). The setpoint for the entire lab was modified? What should the setpoint even be?

Quote:
 

- I went to the south arm. There are two big vent ducts for the outlets and intakes. Both are not flowing the air.
  The current temp at 7pm was ~30degC. Max and min were 31degC and 18degC.

- Then I went to the vertex and the east arm. The outlets and intakes are flowing.

Attachment 1: rmTemp.pdf  623 kB  Uploaded Wed Jul 14 17:21:06 2021  | Hide | Hide all
rmTemp.pdf
ELOG V3.1.3-